Ammonium sulfate reacts with potassium hydroxide solution to produce ammonia gas, potassium sulfate, and water. Calculate the volume of ammonia gas, measured at 23oC and 64 kPa, that could be produced from 264.0 g of ammonium sulfate and 280.0 g of potassium hydroxide
1
Expert's answer
2020-07-27T05:36:28-0400

The volume of ammonia gas produced can be obtained from its number of moles at STP

moles of ammonium sulfate = mass/molar mass

=264/132= 2 moles

(NH4)2(SO4) + 2KOH → K2SO4 + 2NH3 + 2H2O

reacting mole ratio of ammonium sulfate and ammonia gas = 1:2

thus, moles of ammonia gas = 2*2 = 4 moles

pressure of ammonia = 64kPa

temperature = 296K

lets consider STP where volume = 22.4L

pressure = 101.325kPa

temperature = 273K

at STP, the number of moles = 1 mole

if 1 mole, the volume of a gas = 22.4L

at 4 moles, volume = 4*22.4L = 89.6L

using the formula P1V1/T1=P2V2/T2

(101.325*89.6)/273=(64*V2)/296

V2= 153.806L